### v1

# A Gross-Pitaevskii theory for an excitonic incompressible Bose solid

Abstract
We show that interlayer excitons in double-layer semiconductor heterostructures can form a Bose solid which is an incompressible supersolid characterized by exactly one boson per lattice site. This exciton Bose solid would be the first realization of an incompressible supersolid, unlike the generally compressible cluster supersolids seen in dipolar quantum gases. Capturing its characteristics and associated emergent phenomena requires extending the Gross-Pitaevskii formalism to include strong two-particle correlations and exclude exciton self-interactions. We develop such a formalism, we apply it across experimentally accessible exciton densities and interlayer separations, and we show that it incorporates both superfluid and incompressible supersolid ground states. This extended framework allows us to determine the superfluid–supersolid transition and explore the low-temperature properties of the exciton supersolid across its complete parameter space.
For many years there has been considerable theoretical interest in the Bose solid, but it has never been definitively observed experimentally. A true Bose solid must have a lattice site occupancy of exactly one boson per site [1]. Precisely one boson per site generates a particle-antiparticle gap in the energy spectrum, and as a consequence, such a Bose solid will be incompressible. References [2, 3] showed that the ground state of a Bose solid is an incompressible supersolid with the particles localized on their own lattice sites by strong particle-particle interactions, while at the same time they are in a single quantum condensate. The Bose solid/incompressible supersolid is quite different from the superfluids with periodic clustering [4, 5] or density modulation [6] in dipolar cold atom systems that have been termed supersolid-like.
In this letter, we generalize the Gross-Pitaevskii approach to map out the ground state of the Bose solid at zero temperature. The Gross-Pitaevskii formalism is mean-field Hartree, so it includes interactions of a particle with itself [7, 8]. However in a Bose solid with only a single particle on each site, there should be no self-interaction terms [9]. Here, we extend the conventional Gross-Pitaevskii approach to capture this property. Furthermore, exchanges of bosons in the Bose solid may take place with difficulty because of the presence of other neighbors [1]. We therefore also incorporate in our Gross-Pitaevskii approach the effect of strong short-range two-particle correlations.
Our results show that excitons in a double-layer semiconductor heterostructure can form a Bose solid and would constitute the first experimental realization of an incompressible supersolid. In this system, the excitons are spatially indirect, formed of electrons and holes that are confined in two adjacent two-dimensional conducting layers separated by a thin insulating barrier [10, 11]. With equal electron and hole densities, the exciton dipoles will be aligned perpendicular to the layers so the dipole-like exciton interactions are purely repulsive. Sufficiently strong repulsion can localize the excitons on separate lattice sites of a dipolar crystal [12, 13] similarly to electrons in a Wigner crystal [14]. This state differs from cold atom systems where attractive as well as repulsive atomic interactions acting in competition are needed to generate periodic clusters [15] which in general are not incompressible. We have proposed the existence of an exciton incompressible supersolid, based on a variational calculation [16].
The Hamiltonian for the excitonic state is
$\widehat{H}$ $=$ $\int d^{2}\mathbf{r}\,\widehat{\psi}^{\dagger}(\mathbf{r})\left(-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\right)\widehat{\psi}(\mathbf{r})$
$+$ $\frac{1}{2}\!\!\iint d^{2}\mathbf{r}d^{2}\mathbf{r}^{\prime}\,\widehat{\psi}^{\dagger}(\mathbf{r})\widehat{\psi}^{\dagger}(\mathbf{r}^{\prime})V_{\text{XX}}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\widehat{\psi}(\mathbf{r})\widehat{\psi}(\mathbf{r}^{\prime}),$
where $\widehat{\psi}(\mathbf{r})$ ($\widehat{\psi}^{\dagger}(\mathbf{r})$) is the exciton boson destruction (creation) operator at position $\mathbf{r}$ for the order parameter of the quantum condensate, and $M_{\text{X}}=m_{e}^{*}+m_{h}^{*}$ is the exciton mass. We set the electron and hole masses equal, $m_{e}^{*}=m_{h}^{*}$.
We take equal electron and hole densities, the exciton dipoles align perpendicular to the layers, and so the exciton-exciton interactions are purely repulsive,
$V_{\text{XX}}(r)=\frac{2e^{2}}{4\pi\epsilon}\left(\frac{1}{r}-\frac{1}{\sqrt{r^{2}+d^{2}}}\right)\ .$ (2)
Here $d$ is the distance between the layers, $\epsilon$ is the dielectric constant of the insulating layer that separates the layers, and $\mathbf{r}$ is the separation of the excitons in the plane of the layers. $V_{\text{XX}}(r)$ is the sum of the four bare Coulomb interactions that act between the electrons and holes forming the two excitons.
At low densities, the $V_{\text{XX}}$ is well approximated by dipolar repulsion. From Quantum Monte Carlo (QMC) calculations, the pair correlation function $g(\mathbf{r})$ for dipoles at low densities vanishes for small $r$ [12]. This effect is not considered in the conventional Hartree-Fock approach, but we may include it with an effective Hamiltonian ${\cal{H}}^{\rm{eff}}$ in which the effective exciton-exciton interaction contains a hard core of radius $R_{HC}$ that can be deduced from the structure factors taken from Ref. [12]. This keeps the excitons outside the region where the $g(\mathbf{r})$ as determined from QMC vanishes. It is equivalent to considering a local t-matrix $t_{\text{XX}}(\mathbf{r})$, representing multiple short-range two-body scattering, that vanishes for $r<R_{HC}$ [17]. The effective Hamiltonian thus reads
^$\widehat{{\cal{H}}}^{\rm{eff}}=\int$ $d^{2}\mathbf{r}$ $\,\widehat{\psi}^{\dagger}(\mathbf{r})\left(-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\right)\widehat{\psi}(\mathbf{r})$
$\!\!+\frac{1}{2}\!\iint\!$ $d^{2}\mathbf{r}$ $\,d^{2}\mathbf{r}^{\prime}\,\widehat{\psi}^{\dagger}(\mathbf{r})\widehat{\psi}^{\dagger}(\mathbf{r}^{\prime})t_{\text{XX}}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\widehat{\psi}(\mathbf{r})\widehat{\psi}(\mathbf{r}^{\prime}),$ (3)
where
$t_{\text{XX}}(r)$ $=$ $\left\{\begin{array}[]{ll}V_{\text{XX}}(r)\ \ \ \mbox{ if }r>R_{HC};\\ 0\ \ \ \ \ \ \ \ \ \ \ \ \ \mbox{otherwise. }\end{array}\right.$ (6)
For sufficiently strong exciton interactions, a normal-state exciton crystal can form, with exactly one exciton per lattice site [12, 13]. This normal-state Bose solid can be described by a many-body wave function consisting of a product of single-particle wave functions $\phi_{i}(\mathbf{r})$, each $\phi_{i}(\mathbf{r})$ centered on a different site $i$ of the periodic lattice [2]. We derive a Hartree equation for this state in the supplemental material (SM) [18], as
$-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\phi_{i}(\mathbf{r})$ $+$ $\int\!\!d^{2}\mathbf{r}^{\prime}\phi_{i}(\mathbf{r})t_{\text{XX}}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\sum_{\begin{subarray}{c}j\end{subarray}}|\phi_{j}(\mathbf{r}^{\prime})|^{2}$ (7)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!-\int\!\!d^{2}\mathbf{r}^{\prime}\phi_{i}(\mathbf{r})t_{\text{XX}}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\phi_{i}(\mathbf{r}^{\prime})|^{2}=\mu\phi_{i}(\mathbf{r}).$
The attractive interaction term in Eq. (21) is generated from the property that, while an exciton on a singly-occupied site $i$ feels the repulsive interactions from the excitons in the neighboring unit cells, it does not feel its own potential within the unit cell $i$ [1] (see Eq. (8) in the SM [18]). This term removes the contributions in the Hartree equation from a single boson in a unit cell interacting with itself. The term is only present when there is exactly one particle per unit cell, since with more particles there would be interactions within the unit cell. The effect of the absence of these self-interactions within a site acts like a potential well centered on the site. For sufficiently strong repulsion between the excitons, the potential wells can become deep enough for bound-state solutions to exist [1]. In fact, in systems with purely repulsive particle-particle interactions such as electrons or perpendicularly aligned excitons, this is the mechanism that can drive the solidification [2], and in these repulsive systems quantum Monte Carlo simulations have confirmed the existence of normal solids [14, 12, 13]. The solid would be incompressible since the potential wells lead to a prohibitive energy cost against doubly occupied sites [9], analogous to a Mott insulator.
In the present study we investigate the possibility of solidification coexisting with quantum coherence in a supersolid. We extend the standard Gross-Pitaevskii formalism to include the effect of the absence of self interactions, specific to this exciton system. In place of the many-body wave function of the normal state system, we construct the order parameter for the incompressible supersolid of the form, $\Psi(\mathbf{r})=\sum_{i=1}^{N}\phi_{i}(\mathbf{r})$, where $i$ is the site index for the $N$ sites. With this construction, each exciton has the same wave function spread across the entire lattice and with peaks at every site. This order parameter results in both off-diagonal long-range order and also diagonal long-range order coming from the lattice symmetry. In addition, for a supersolid there must exist phase locking between adjacent sites. This arises from mutual overlap of the $\phi_{i}(\mathbf{r})$ [19]. If the $\phi_{i}(\mathbf{r})$ were exponentially localized, it becomes impossible to link the phases on the different lattice sites, since thermal and zero-point quantum fluctuations would eliminate any coherence between the near-negligible overlaps. In order to monitor the overlaps, we introduce a parametrization of the order parameter. We represent $\phi_{i}$ and $\Psi$ in the form,
$\phi_{i}(\mathbf{r})$ $=$ $\frac{\sqrt{\rho_{B}}}{N}+G(\mathbf{r}-\mathbf{r}_{i})\ ,$ (8)
$\Psi(\mathbf{r})$ $=$ $\frac{1}{\sqrt{N}}\left(\sqrt{\rho_{B}}+\sum_{i=1}^{N}G(\mathbf{r}-\mathbf{r}_{i})\right)\ .$ (9)
This measures the overlap by means of a uniform background component $\rho_{B}$. In cases when $\rho_{B}$ tends to zero there is negligible overlap so there would be no supersolid. The $G(\mathbf{r}-\mathbf{r}_{i})$ are functions localized on the different sites $i$ and centered at $\mathbf{r}_{i}$. In cases when the $G(\mathbf{r}-\mathbf{r}_{i})$ reduce to zero, Eq. (24) represents the homogeneous order parameter for a superfluid, $|\Psi(\mathbf{r})|^{2}\sim\rho_{B}$.
From Eq. (21), we derive a time-independent Gross-Pitaevskii equation [20] for the order parameter $\Psi(\mathbf{r})$ (see the SM [18]),
$\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\Psi(\mathbf{r})+U_{\text{XX}}(\mathbf{r})\Psi(\mathbf{r})=\mu\Psi(\mathbf{r})\ ,$ (10)
with mean-field potential,
$U_{\text{XX}}(\mathbf{r})$ $=$ $\int d^{2}\mathbf{r}^{\prime}\,t_{\text{XX}}(|\mathbf{r}^{\prime}-\mathbf{r}|)|\Psi(\mathbf{r}^{\prime})|^{2}$ (11)
$-\int_{\mathbf{r}^{\prime}\in i_{\mathbf{r}}}\!\!\!\!\!\!d^{2}\mathbf{r}^{\prime}t_{\text{XX}}(|\mathbf{r}^{\prime}-\mathbf{r}|)\,|\Psi(\mathbf{r}^{\prime})|^{2}\ .$
The notation ${\mathbf{r}^{\prime}\in i_{\mathbf{r}}}$ indicates that the integral over $\mathbf{r}^{\prime}$ is restricted to the same $i_{\mathbf{r}}$ unit cell as $\mathbf{r}$.
The potential $U_{XX}(\mathbf{r})$ differs from the potential in the standard Gross-Pitaevskii equation for dipolar bosons [21], not only because $V_{\text{XX}}$ has been replaced by $t_{\text{XX}}$ to account for two-body correlations, but also because of the additional potential term. It is this term, that accounts for the absence of self-interactions, which makes supersolid solutions possible.
Figure 1: Profile along diagonal and the spatial distribution of order parameter $|\Psi(\mathbf{r})|^{2}$ for $r_{0}=30$. Panels (a)–(d) correspond to the interlayer separations $d=1,7,15,20$, respectively. Shaded area shows the homogeneous background component of $|\Psi(\mathbf{r})|^{2}$.
We solve Eqs. (10) and (11) by evolving in imaginary time [22], starting from a spatially randomized order parameter. Note throughout the numerical calculation there is no restriction that $\Psi(\mathbf{r})$ is of the form of Eq. (24). Because of the non-local interaction $V_{\text{XX}}(\mathbf{r})$, we evaluate Eq. (11) by means of the convolution theorem [23]. For simplicity, we define the $i_{\mathbf{r}}$ in Eq. (11) on a square lattice. For exciton density $\rho$, the lattice constant is given by $a=2/(\pi\sqrt{\rho})$ for one exciton per site. A triangular lattice may have a slightly lower energy, but we leave the issue of square vs. triangular lattice outside the scope of the present study. All simulations were performed on a $256^{2}$ grid covering the sample spatial area. Solutions were taken as converged when the relative change in energy in consecutive time steps decreases below $10^{-8}$. After reaching convergence, we have verified for self-consistency that $\Psi(\mathbf{r})$ indeed has the form of Eq. (24), with $\rho_{B}$ not zero.
We focus first on the low-density regime, setting the average interparticle spacing $r_{0}=1/(\pi\sqrt{\rho})=30$. As length and energy scales, we use the effective Bohr radius $a_{B}=\hbar^{2}\epsilon/e^{2}m_{e}^{*}$ and the effective Rydberg Ry. Figure 1(a) shows the ground-state order parameter $|\Psi(\mathbf{r})|^{2}$ for small interlayer separation $d=1$. The solution $\Psi(\mathbf{r})$ is homogeneous – corresponding to a superfluid: for such a small $d$, the exciton dipole interactions are too weak to drive the system to a solid.
It is interesting to note that for this small value of $d$, the solution is unchanged if in Eq. (11) we do not include the negative term that removes the self-interaction. The self-interactions thus have no effect on the ground state for such weak dipolar interactions.
By increasing $d$ we increase the dipole moment, making the exciton interactions stronger. In Fig. 1(b) for $d=7$, we see that the order parameter $\Psi(\mathbf{r})$ is now inhomogeneous with a periodicity of the lattice. The $\Psi(\mathbf{r})$ has a normalized peak at each lattice site and there is a large overlap between adjacent sites. Thus we conclude that this is an incompressible supersolid ground state, with the order parameter $\Psi(\mathbf{r})$ indeed of the parametrized form of Eq. (24). The overlap that allows exchange of bosons is associated with the homogeneous background component of amplitude $\rho_{B}$ (shaded area).
Figures 1(c) and 1(d) show that when the interlayer spacing $d$ is further increased, the peak in $\Psi(\mathbf{r})$ on each site narrows and the amplitude $\rho_{B}$ of the homogeneous background component decreases.
Figure 2(a) shows the dependence of the amplitude $\rho_{B}$ on $d$ for fixed $r_{0}=30$. The behavior of $\rho_{B}$ is similar to that obtained in the variational calculation of Ref. [16]. $\rho_{B}/\rho=1$ corresponds to a homogeneous superfluid solution, while for a supersolid solution, $0<\rho_{B}/\rho<1$. The background component decreases with increasing $d$ but for a supersolid solution it does not vanish.
Figure 2(b) compares the supersolid energy eigenvalue $\mu_{SS}$ from Eq. (10) determined using $U_{XX}(\mathbf{r})$ from Eq. (11) and the superfluid energy eigenvalue $\mu_{SF}$ determined without the subtracted self-interaction term in $U_{XX}(\mathbf{r})$. For $d<6$ the two energy eigenvalues are identical, indicating a stable superfluid, but for $d\gtrsim 6$ the energy eigenvalue $\mu_{SF}$ is higher than $\mu_{SS}$ indicating that the incompressible supersolid becomes the ground state for $d\gtrsim 6$.
Repeating the above analysis for increasing density allows us to map out the $r_{0}$-$d$ phase diagram at zero temperature, shown in Figure 3. For a fixed low density ($r_{0}=30$) and small $d$, the exciton interaction $V_{\text{XX}}$ is weak, leading to a superfluid coherent phase with a constant order parameter $\Psi(\mathbf{r})$ (inset at $d=2.5$). When $d$ is increased, our Gross-Pitaevskii formalism predicts a transition from the superfluid to an incompressible supersolid (dashed red line). This lies very close to the corresponding transition line predicted from the variational calculation in Ref. [16].
At very large $d$ a normal exciton solid is expected to become the ground state [12, 13, 16]. In contrast to the supersolid, in the normal exciton solid, the peaks of the wave function are independently localized without exchange. Interestingly, our Gross-Pitaevskii approach does reflect an evolution towards the normal exciton solid for increasing $d$: the peaks in the order parameter become narrower and the amplitude of the homogeneous background trends towards zero (insets for $d=30$). However, the Gross-Pitaevskii formalism cannot determine the transition to the normal exciton solid since it a priori assumes a global phase coherence of one order parameter with a non-zero homogeneous background.
Figure 2: (a) Dependence on interlayer separation $d$ of $\rho_{B}/\rho$, the amplitude of the homogeneous background component of $\Psi(\mathbf{r})$, for $r_{0}=30$. (b) Energy eigenvalue of the supersolid state (SS) and corresponding eigenvalue of the superfluid state (SF), calculated without the self-interaction correction.
Figure 3 also shows the effect on the supersolid $\Psi(\mathbf{r})$ of increasing the density for a fixed $d$ (see insets at $d=15$). As $r_{0}$ decreases, the peaks in $\Psi(\mathbf{r})$ become broader and the homogeneous background component of $\Psi(\mathbf{r})$ grows larger. This is associated with the evolution with $r_{0}$ of the average exciton interaction $V_{\text{XX}}$ (Eq. (2)) from a dipolar form to a coulombic form. Coulomb interactions are known to support crystallization only at very low densities [24, 25], and so decreasing $r_{0}$ has the effect of softening the supersolid. The supersolid phase shown in Fig. 3 lies almost everywhere in the BEC regime, with condensate fractions of order unity, where screening is negligible or very weak [26]. However, before the supersolid melts with decreasing $r_{0}$, the exciton condensate passes into the BEC-BCS crossover regime where screening would rapidly switch on [27], killing the quantum coherence. The transition from quantum coherence to the normal state is indicated by the grey dash-dotted line [28, 29]. In this high-density region of phase space, neither coulombic nor dipolar interactions support crystallization in the normal state [24, 25, 12]. Thus we expect that when the quantum coherence is lost, simultaneously the supersolid will melt to the normal-state liquid.
Since the BEC to BEC-BCS crossover boundary lies so close to the melting curve [28], our Gross-Pitaevskii formalism with bosonic single particles interacting without screening should be valid for essentially all of the area covered by the superfluid and supersolid phases shown in Fig. 3.
In conclusion, we have demonstrated that an incompressible Bose solid/supersolid with one particle per site [1] should be experimentally accessible for a system of excitons in an electron-hole double-layer heterostructure. Our approach is based on the Gross-Pitaevskii formalism and can map out the quantum condensed phases over the zero-temperature phase diagram. Starting with the formalism for aligned bosonic excitons with purely repulsive dipolar-like interactions, we extended the formalism (i) to exclude the self-interaction energies, which should be absent for single occupancy sites [3], and (ii) to take into account strong two-particle correlations.
For weak exciton-exciton interactions our formalism predicts a superfluid ground state, but for stronger interactions there is a transition to an incompressible supersolid/Bose solid, with an order parameter that has peaks on a periodic lattice of sites superimposed on a uniform background. The background component found in our calculations is necessary for the phase coherence of the incompressible supersolid [19]. For even stronger interactions, our formalism displays an evolution towards a normal exciton solid, with the peaks in the order parameter narrowing and increasing in amplitude while its uniform background component decreases towards zero. In a normal exciton solid, the peaks would be independently localized without phase coherence between the peaks. In that respect, the evolution from supersolid to normal solid resembles that observed in the superfluid to Mott insulator transition in quantum gases [30].
Figure 3: The phase diagram at zero temperature, as a function of interlayer separation $d$ and average interparticle spacing $r_{0}$. Dashed red line: superfluid to supersolid transition. Dash-dotted grey line: condensate collapse to a normal liquid [29]. Dashed green line: exciton liquid to normal solid transition from Ref. [12]. Dotted black line: $d=r_{0}$. Insets: order parameter profile $|\Psi(\mathbf{r})|^{2}$ at the points indicated by grey dots.
Within the supersolid phase, as the density is increased the exciton repulsive interaction (Eq. (2)) evolves from dipolar-like to Coulomb-like, making the supersolid soften and eventually melt. Just before melting, the system enters the BEC-BCS crossover regime where screening rapidly becomes strong and the system simultaneously loses coherence and crystal structure. The quantum condensed phases lie almost entirely within the BEC regime where screening is negligible, allowing us to describe the system in terms of a Bose condensate with a macroscopic order parameter. This makes our Gross-Pitaevskii formalism valid across the area of the phase diagram encompassing the supersolid and superfluid phases.
Our Gross-Pitaevskii formalism establishes strong foundations for exploring the elusive incompressible supersolid and for identifying its hallmark features. For example, by studying the formation of vortices and vortex lattices [31, 32, 33] and extending the formalism to the time-dependent Gross-Pitaevskii equation, the dynamics of the vortices [34, 35] can be used to identify the transition to an incompressible supersolid. This powerful framework also gives access to the rich dynamics of phase transitions to investigate critical phenomena and scaling behavior, as well as access to the collective excitations of the condensate, capturing the behavior of the low-energy collective modes to provide insights into the thermodynamics of the Bose solid.
Acknowledgements
We thank Filippo Pascucci for very useful discussions. This work was supported by Research Foundation – Flanders (FWO-Vl), with contract number 12A3T24N; by the University Research Fund (BOF), with project numbers 53228, G0A9F25N, G0AIY25N, GOH1122N, G061820N, and G060820N; and by the Brazilian Council for Research (CNPq), through the PRONEX/FUNCAP, Universal, and PQ programs.
Appendix A Derivation of Gross-Pitaevskii equation for an exciton Bose solid
We outline our generalized Gross-Pitaevskii formalism. Consider the wave function for a Bose solid [2] in the normal state,
|$\ket{\Psi_{\mathrm{BS}}(\mathbf{r})}=\displaystyle\prod_{m=1}^{N}\hat{c}_{m}^{\dagger}\ket{\Psi_{\mathrm{vac}}(\mathbf{r})}\ .$ (12)
|$\ket{\Psi_{\mathrm{vac}}(\mathbf{r})}$ is the vacuum, $\hat{c}_{m}^{\dagger}=\int d^{2}\mathbf{r}\,\phi_{m}^{\star}(\mathbf{r})\widehat{\psi}^{\dagger}(\mathbf{r})$, $\hat{c}_{m}=\int d^{2}\mathbf{r}\,\phi_{m}(\mathbf{r})\widehat{\psi}(\mathbf{r})$. The $\phi_{m}(\mathbf{r})$ are a set of identical functions, each centered on the different sites $\{m\}$ of a periodic lattice. They are degenerate and mutually orthogonal.
The effective Hamiltonian (Eq. (333) in the main manuscript) can be written,
^$\widehat{{\cal{H}}}^{\rm{eff}}=\sum_{i,j}\hat{c}_{i}^{\dagger}T_{ij}\hat{c}_{j}+\frac{1}{2}\!\!\sum_{i,i^{\prime},j,j^{\prime}}\hat{c}_{i}^{\dagger}\hat{c}_{i^{\prime}}^{\dagger}W_{i,i^{\prime},j,j^{\prime}}\hat{c}_{j}\hat{c}_{j^{\prime}}\ ,$ (13)
with kinetic and potential energy ground state expectation values,
$T_{ij}$ $=$ $\int\,d^{2}\mathbf{r}\,\phi_{i}^{\star}(\mathbf{r})\left(-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\right)\phi_{j}(\mathbf{r})\ ,$ (14)
$W_{i,i^{\prime},j,j^{\prime}}\!\!$ $=$ $\!\!\iint\!\!d^{2}\mathbf{r}\,d^{2}\mathbf{r}^{\prime}\phi_{i}^{\star}(\mathbf{r})\phi_{i^{\prime}}^{\star}(\mathbf{r}^{\prime})t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\phi_{j}(\mathbf{r}^{\prime})\phi_{j^{\prime}}(\mathbf{r})\ .$ (15)
The expectation energy functional for the state |$\ket{\Psi_{\mathrm{BS}}(\mathbf{r})}$ is,
$\bra{\Psi_{\mathrm{BS}}(\mathbf{r})}\widehat{{\cal{H}}}^{\rm{eff}}\ket{\Psi_{\mathrm{BS}}(\mathbf{r})}=$ $\left(\bra{{\psi}_{\mathrm{vac}}(\mathbf{r})}\displaystyle\prod_{m=1}^{N}\hat{c}_{m}\right)\sum_{i,j}\hat{c}_{i}^{\dagger}T_{ij}\hat{c}_{j}\left(\displaystyle\prod_{m^{\prime}=1}^{N}\hat{c}_{m^{\prime}}^{\dagger}\ket{{\psi}_{\mathrm{vac}}(\mathbf{r})}\right)$ (16)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\left(\bra{{\psi}_{\mathrm{vac}}(\mathbf{r})}\displaystyle\prod_{m=1}^{N}\hat{c}_{m}\right)\frac{1}{2}\!\!\sum_{\begin{subarray}{c}i,i^{\prime},j,j^{\prime}\end{subarray}}\hat{c}_{i}^{\dagger}\hat{c}_{i^{\prime}}^{\dagger}W_{i,i^{\prime},j,j^{\prime}}\hat{c}_{j}\hat{c}_{j^{\prime}}\left(\displaystyle\prod_{m^{\prime}=1}^{N}\hat{c}_{m^{\prime}}^{\dagger}\ket{{\psi}_{\mathrm{vac}}(\mathbf{r})}\right)\ .$
Note that in the interaction term of Eq. (16) the contributions in the summation $j\!=\!j^{\prime}$ and $i\!=\!i^{\prime}$ will give zero. This is because for the state |$\ket{\Psi_{\mathrm{BS}}}$ there is only one particle on each site $j$ , so $\left(\hat{c}_{j}\right)^{2}\ket{\Psi_{\mathrm{BS}}}=0$ and $\bra{\Psi_{\mathrm{BS}}}\left(\hat{c}_{j}^{\dagger}\right)^{2}\!=0$. We will see that this property is crucial for obtaining the boson solid. Then,
$\bra{\Psi_{\mathrm{BS}}(\mathbf{r})}\widehat{{\cal{H}}}^{\rm{eff}}\ket{\Psi_{\mathrm{BS}}(\mathbf{r})}$
$=\sum_{i,j}\left(\bra{{\psi}_{\mathrm{vac}}(\mathbf{r})}\displaystyle\prod_{\begin{subarray}{c}m=1\\ m\neq i\end{subarray}}^{N}\hat{c}_{m}\right)T_{ij}\left(\displaystyle\prod_{\begin{subarray}{c}m^{\prime}=1\\ m^{\prime}\neq j\end{subarray}}^{N}\hat{c}_{m^{\prime}}^{\dagger}\ket{{\psi}_{\mathrm{vac}}(\mathbf{r})}\right)+\!\frac{1}{2}\!\!\sum_{\begin{subarray}{c}i,i^{\prime},j,j^{\prime}\\ i\neq i^{\prime},j\neq j^{\prime}\end{subarray}}\left(\bra{{\psi}_{\mathrm{vac}}(\mathbf{r})}\displaystyle\prod_{\begin{subarray}{c}m=1\\ m\neq i,i^{\prime}\end{subarray}}^{N}\hat{c}_{m}\right)W_{i,i^{\prime},j,j^{\prime}}\left(\displaystyle\prod_{\begin{subarray}{c}m^{\prime}=1\\ m^{\prime}\neq j,j^{\prime}\end{subarray}}^{N}\hat{c}_{m^{\prime}}^{\dagger}\ket{{\psi}_{\mathrm{vac}}(\mathbf{r})}\right)$
$=\sum_{i,j}T_{ij}\delta_{ij}+\frac{1}{2}\!\!\sum_{\begin{subarray}{c}i,i^{\prime},j,j^{\prime}\\ i\neq i^{\prime},j\neq j^{\prime}\end{subarray}}W_{i,i^{\prime},j,j^{\prime}}\left(\delta_{ij}\delta_{i^{\prime}j^{\prime}}+\delta_{ij^{\prime}}\delta_{i^{\prime}j}\right)$
$=\sum_{i}\int d^{2}\mathbf{r}\,\phi_{i}^{\star}(\mathbf{r})\left(-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\right)\phi_{i}(\mathbf{r})$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}\\ j\neq j^{\prime}\end{subarray}}\iint d^{2}\mathbf{r}\,d^{2}\mathbf{r}^{\prime}\left(\phi_{j}^{\star}(\mathbf{r})\phi_{j^{\prime}}^{\star}(\mathbf{r}^{\prime})t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\phi_{j}(\mathbf{r}^{\prime})\phi_{j^{\prime}}(\mathbf{r})+\phi_{j^{\prime}}^{\star}(\mathbf{r})\phi_{j}^{\star}(\mathbf{r}^{\prime})t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\phi_{j}(\mathbf{r}^{\prime})\phi_{j^{\prime}}(\mathbf{r})\right)$ (17)
Using completeness $\sum_{j}\phi_{j}^{\star}(\mathbf{r})\phi_{j}(\mathbf{r}^{\prime})\!=\!\delta(\mathbf{r}-\mathbf{r}^{\prime})$, we recognize that the first term in the interaction in Eq. (17) is zero since $t_{\text{XX}}(|\mathbf{r^{\prime}}-\mathbf{r}|\!=\!0)=0$, leaving,
$\bra{\Psi_{\mathrm{BS}}(\mathbf{r})}\widehat{{\cal{H}}}^{\rm{eff}}\ket{\Psi_{\mathrm{BS}}(\mathbf{r})}=\sum_{i}\int\,d^{2}\mathbf{r}\,\phi_{i}^{\star}(\mathbf{r})\left(-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\right)\phi_{i}(\mathbf{r})+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}\\ j\neq j^{\prime}\end{subarray}}\iint\!\!d^{2}\mathbf{r}\,d^{2}\mathbf{r}^{\prime}|\phi_{j^{\prime}}(\mathbf{r})|^{2}t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\phi_{j}(\mathbf{r}^{\prime})|^{2}\ .$ (18)
The term $j=j^{\prime}$ that is absent from the sum in Eq. (18) can alternatively be accounted for by the subtraction of a potential from an unrestricted sum over $i$ and $j$. Then Eq. (16) becomes,
$E(\phi_{i},\phi_{i}^{*})=\sum_{i}\int\,d^{2}\mathbf{r}\,\phi_{i}^{\star}(\mathbf{r})\left(-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\right)\phi_{i}(\mathbf{r})$ $+$ $\frac{1}{2}\sum_{\begin{subarray}{c}i,j\end{subarray}}\iint\!\!d^{2}\mathbf{r}\,d^{2}\mathbf{r}^{\prime}|\phi_{i}(\mathbf{r})|^{2}t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\phi_{j}(\mathbf{r}^{\prime})|^{2}$ (19)
-- $\frac{1}{2}\sum_{i}\iint\!\!d^{2}\mathbf{r}\,d^{2}\mathbf{r}^{\prime}|\phi_{i}(\mathbf{r})|^{2}t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\phi_{i}(\mathbf{r}^{\prime})|^{2}\ .$
We make use of a variational approach to obtain the ground state solution [20], Eq. (19) is minimized under variations $\delta\phi_{i}^{\star}(\mathbf{r})$. A Lagrange multiplier $\mu$ ensures the total number of particles is conserved, $\int d^{2}\mathbf{r}|\phi_{i}(\mathbf{r})|^{2}=1$,
$\frac{\delta}{\delta\phi_{i}^{*}}\left(E(\phi_{i},\phi_{i}^{*})-\mu\sum_{i}\int d^{2}\mathbf{r}|\phi_{i}(\mathbf{r})|^{2}\right)=0\ .$ (20)
Equation (20) gives,
$-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\phi_{i}(\mathbf{r})$ $+$ $\int\!\!d^{2}\mathbf{r}^{\prime}\phi_{i}(\mathbf{r})t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\sum_{\begin{subarray}{c}j\end{subarray}}|\phi_{j}(\mathbf{r}^{\prime})|^{2}-\int\!\!d^{2}\mathbf{r}^{\prime}\phi_{i}(\mathbf{r})t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\phi_{i}(\mathbf{r}^{\prime})|^{2}=\mu\phi_{i}(\mathbf{r})\ .$ (21)
For an exciton quantum condensate, the macroscopic occupation of the condensate implies that a single-particle function, the order parameter, describes the many-body ground state. Instead of the many-body wave function for the normal state system (Eq. 12 that consists of a product of different single-particle wave functions each to describe an exciton on its site) the order parameter that we choose for the incompressible supersolid is of the form,
$\Psi(\mathbf{r})$ $=$ $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\phi_{i}(\mathbf{r})\ ,$ (22)
for $N$ sites with $N$ particles. A system described by the order parameter ((22)) has off-diagonal long-range order in the presence of a non-zero overlap between adjacent sites, establishing phase coherence. The $\Psi(\mathbf{r})$ will have a maximum at every lattice site $i$, so additionally it also has diagonal long-range order due to the lattice symmetry. The combination of both off-diagonal and diagonal long-range order makes Eq. (22) a candidate for a supersolid order parameter.
For a supersolid there must exist phase locking between adjacent sites arising from mutual overlap of the $\phi_{i}(\mathbf{r})$ [19]. To monitor the overlap, we parametrize the order parameter, writing $\phi_{i}$ in the form,
$\phi_{i}(\mathbf{r})$ $=$ $\frac{\sqrt{\rho_{B}}}{N}+G(\mathbf{r}-\mathbf{r}_{i})$ (23)
$\Psi(\mathbf{r})$ $=$ $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\left(\frac{\sqrt{\rho_{B}}}{N}+G(\mathbf{r}-\mathbf{r}_{i})\right)\ .$ (24)
The background component $\sqrt{\rho_{B}}$ is uniform throughout the sample, and all the $G(\mathbf{r}-\mathbf{r}_{i})$ are the same function but localized on the different sites $i$ centered at $\mathbf{r}_{i}$.
Summing Eq. (21) over the sites $i$ and using Eq. (22) we obtain,
$-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\sum_{i}\phi_{i}(\mathbf{r})+\sum_{i}\!\!\int\!\!d^{2}\mathbf{r}^{\prime}\phi_{i}(\mathbf{r})t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\sum_{\begin{subarray}{c}j\end{subarray}}|\phi_{j}(\mathbf{r}^{\prime})|^{2}-\sum_{i}\!\int\!\!d^{2}\mathbf{r}^{\prime}\phi_{i}(\mathbf{r})t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\phi_{i}(\mathbf{r}^{\prime})|^{2}=\mu\!\sum_{i}\phi_{i}(\mathbf{r})$ (25)
$-\sqrt{N}\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\Psi(\mathbf{r})+\sqrt{N}\!\int\!\!d^{2}\mathbf{r}^{\prime}\Psi(\mathbf{r})\ t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\sum_{\begin{subarray}{c}j\end{subarray}}|\phi_{j}(\mathbf{r}^{\prime})|^{2}-\sum_{i}\!\!\int\!\!d^{2}\mathbf{r}^{\prime}\phi_{i}(\mathbf{r})\ t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\phi_{i}(\mathbf{r}^{\prime})|^{2}=\sqrt{N}\mu\Psi(\mathbf{r})$ (26)
From Eq. (23), within the second term of Eq. (26) we can write,
$\sum_{j}|\phi_{j}(\mathbf{r})|^{2}$ $=$ $\sum_{j}\left|\frac{\sqrt{\rho_{B}}}{N}+G(\mathbf{r}-\mathbf{r}_{j})\right|^{2}=\left|\frac{\rho_{B}}{N}+2\frac{\sqrt{\rho_{B}}}{N}\sum_{j}G(\mathbf{r}-\mathbf{r}_{j})+\sum_{j}G^{2}(\mathbf{r}-\mathbf{r}_{j})\right|\ ,$ (27)
When $G(\mathbf{r}-\mathbf{r}_{i})$ is localized within a single unit cell $i$, then for a value of $\mathbf{r}^{\prime}=\mathbf{r}_{0}$ located in the unit cell $i_{0}$, the summations over $j$ collapse to,
$\sum_{i}|\phi_{i}(\mathbf{r}_{0})|^{2}$ $=$ $\frac{\rho_{B}}{N}+2\frac{\sqrt{\rho_{B}}}{N}G(\mathbf{r}_{0}-\mathbf{r}_{i_{0}})+G^{2}(\mathbf{r}_{0}-\mathbf{r}_{i_{0}})\ .$ (28)
Using Eq. (24),
$|\Psi(\mathbf{r}_{0})|^{2}$ $=$ $\frac{\rho_{B}}{N}+2\frac{\sqrt{\rho_{B}}}{N}\sum_{i}G(\mathbf{r}_{0}-\mathbf{r}_{i})+\frac{1}{N}\sum_{i}G(\mathbf{r}_{0}-\mathbf{r}_{i})\sum_{i^{\prime}}G(\mathbf{r}-\mathbf{r}_{i^{\prime}})\ .$ (29)
The final term $\sum_{i}G(\mathbf{r}_{0}-\mathbf{r}_{i})\sum_{i^{\prime}}G(\mathbf{r}_{0}-\mathbf{r}_{i^{\prime}})$ in Eq. (29) is periodic since it is invariant with shifts in $\mathbf{r}\rightarrow\mathbf{r}+\mathbf{n}\cdot\mathbf{a}$, for $\mathbf{n}=\{n_{x},n_{y}\}$ any pair of integers. So in the summation over $i$, we may choose for each $i$ a different value for $\mathbf{n}_{i}$ such that $(\mathbf{r}_{i}+\mathbf{n}_{i}\cdot\mathbf{a})$ is always the unit cell containing $\mathbf{r}_{0}$. Using this construction, each term in the summation over $i$ becomes equal. Thus we may truncate the summation over $i$ to the single value $i=i_{0}$, multiplying by a factor $N$ to account for the $N$ equal terms,
$\sum_{i}G(\mathbf{r}_{0}-\mathbf{r}_{i})\sum_{i^{\prime}}G(\mathbf{r}_{0}-\mathbf{r}_{i^{\prime}})$ $=$ $\sum_{i}\left\{G(\mathbf{r}_{0}-\mathbf{r}_{i}-\mathbf{n}_{i}\cdot\mathbf{a})\sum_{i^{\prime}}G(\mathbf{r}_{0}-\mathbf{r}_{i^{\prime}}-\mathbf{n}_{i}\cdot\mathbf{a})\right\}$ (30)
$=$ $NG(\mathbf{r}_{0}-\mathbf{r}_{i_{0}})\sum_{i^{\prime}}G(\mathbf{r}_{0}-\mathbf{r}_{i^{\prime}}-\mathbf{n}_{i_{0}}\cdot\mathbf{a})$
$=$ $NG^{2}(\mathbf{r}_{0}-\mathbf{r}_{i_{0}})\ ,$ (31)
since in the sum over $i^{\prime}$ in Eq. (30), only the term $(\mathbf{r}_{i^{\prime}}+\mathbf{n}_{i_{0}}\cdot\mathbf{a})=\mathbf{r}_{i_{0}}$ is non-zero. Equation (29) can thus be written,
$|\Psi(\mathbf{r}_{0})|^{2}$ $=$ $\frac{\rho_{B}}{N}+2\frac{\sqrt{\rho_{B}}}{N}G(\mathbf{r}_{0}-\mathbf{r}_{i_{0}})+\frac{1}{N}NG^{2}(\mathbf{r}_{0}-\mathbf{r}_{i_{0}})\ ,$ (32)
and we conclude from Eqs. (28) and (32),
$|\Psi(\mathbf{r})|^{2}=\sum_{i}|\phi_{i}(\mathbf{r})|^{2}\ ,$ (33)
Hence in Eq. (26) the second term becomes,
$\sqrt{N}\!\int\!\!d^{2}\mathbf{r}^{\prime}\Psi(\mathbf{r})\ t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\sum_{\begin{subarray}{c}j\end{subarray}}|\psi_{j}(\mathbf{r}^{\prime})|^{2}=\sqrt{N}\!\int\!\!d^{2}\mathbf{r}^{\prime}\Psi(\mathbf{r})\ t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\Psi(\mathbf{r}^{\prime})|^{2}\ .$ (34)
In Eq. (26), the final term on the left-hand side has the same structure as the second term, but with the restriction $j=i$. We can accordingly use the same construction, but we must capture the index restriction $j=i$ by limiting the $\mathbf{r}^{\prime}$ integration to the unit cell containing $\mathbf{r}$. We use the notation ${\mathbf{r^{\prime}}\in{j_{\mathbf{r}}}}$. So we finally obtain,
$-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\Psi(\mathbf{r})$ $+$ $\int\!\!d^{2}\mathbf{r}^{\prime}\Psi(\mathbf{r})\ t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\Psi(\mathbf{r}^{\prime})|^{2}-\int_{\mathbf{r^{\prime}}\in{j_{\mathbf{r}}}}\!\!d^{2}\mathbf{r}^{\prime}\Psi(\mathbf{r})\ t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\Psi(\mathbf{r}^{\prime})|^{2}=\mu\Psi(\mathbf{r})$ (35)
Equation (35) corresponds to Eqs. (7) and (8) in the main manuscript.